# A projection method for particle resampling

### v6

@@ 7 Refinement studies and long time stability
The choice in timestep is based on previous works with PETSc-PIC [13, 14] and Figure 8(a) shows little difference with a smaller time step than used in these studies.
@@ 8 Side effects of resampling
Figure 4(a) demonstrates that the dynamics of the plasma is not highly sensitive the resampling period within a broad range (eg, $1-8$ $[\omega_{p}^{-1}]$), but that rebound starts a little earlier with higher resampling rates indicating that resampling is disturbing the plasma to some extent. To investigated this effect, Figure 9 (left) shows a linear example with resampling every time step and increasing resampling periods. The highest resampling rates clearly lead to instabilities and these instabilities demonstrate classic growing oscillations up to a small amplitude, see detail in Figure 9 (right), that is growing in mean value and plateaus at about $1.0$. Perhaps stability analysis could provide insight on this issue.
@@
While this instability at high resampling rates is not a practical problem, per se, in that all of our tests are stabilized with much lower resampling frequencies, this data does indicate that resampling can have adverse effects. This test problem is highly idealized and does not provide convincing experimental evidence that the adaptivity strategy of simply mapping make to the original grid will be robust in practice. More sophisticated adaptivity strategies are the subject of future work and would require more scientifically relevant models such as problems in magnetic reconnection in at least $2X+3V$.
@@ 8.1 Discussion
Figure 7 sheds light onto the complex relationship and dependency between all the simulation variables, i.e. resampling rate, spatial and velocity grid density and time step. It is difficult to draw definitive conclusions about what exactly is driving the dynamics of the electric field damping and rebounding. As discussed in §7, the grid resolution plays a large role in reducing the interaction of subgrid modes with the physical waves, while the particle density reduces stochastic, noise-driven heating effects in the plasma. These effects do not, however, converge monotonically suggesting that they are linked in some way. This requires further study in future work.
The inclusion of collisions may also provide a smooth mechanism. The collision operator acts as a controlled, physically motivated diffusion term in the kinetic equations. This diffusive term continuously smooths out any build up of high-frequency components in the distribution function which can lead to spurious rebounds. In other algorithms (Fig. 5.5 [21]), clean decays to machine precision have been shown using collisions. Collisions must be used with some level of caution, however. Much of the essential collisionless interactions in the Landau damping test, in particular in the nonlinear regime, rely on fine-scale interactions between the particles and waves. Introducing a collision operator can have the effect of smoothing out those essential fine-scale interactions which, if too strong, will disrupt the fundamental physics of the system. Recent work in the PETSc library [2, 35, 1, 3] has provided a variety of collision algorithms that can be included in the PETSc-PIC framework in future work. Other methods, such as [47], have reported clean decay without collisions, but the details of the numerical method are not entirely clear and some form of dissipation my be present.
@@ 9 Conclusion
The paper develops a new approach to particle resampling that uses a conservative projection, a pseudoinverse, to map any distribution of particles to essentially any other distribution of particle while conserving all moments up to the degree of polynomial that the projection function space can represent exactly. This method is evaluated with a static particle and continuum Cartesian grids, simply remapping to the original particle grid, on standard Landau damping test problems. Tests where noise is problematic, the linear cases, show that resampling reduces noise considerably and coherent dynamics are maintained for long times, where as the solution becomes essentially all noise without resampling.
@@
With entropy measures, from our particle Landau collision operator [35], we can determine the continuum grids required for resampling or a continuum collision operator [3, 35] to keep entropy generation by the projection well below the entropy generated by the physics.
@@
AMR in velocity and regular particle grids on each cell of an adapted grid, like a cubed sphere [3], is a path for generating adapted particle grids through continuum mesh adaptation.
@@ Appendix A Artifact description and reproducibility
PETSc output files with all data, provenance information, and reproducibility instructions for all tables and plots can be obtained from git@gitlab.com:markadams4/resampling-paper.git. This includes the python scrips that generates the plots and run scripts, makefiles and PETSc resource files used to generate the data, and the test harness code in src. The src/A.X directories has data for $A=0.$X. The exact PETSc versions (SHA1) are in the data files, with the provenance data, all parameters used in each test, but any PETSc version from v3.22 should suffice to reproduce this data.

### v9

@@ 7.1 Landau damping
The choice in timestep is based on previous works with PETSc-PIC [Finn2023, Finn2025] and Figure 11(a) shows little difference with a smaller time step than used in these studies.
7.2 Two-stream
As with the Landau damping tests, for long time stability, two-stream instability runs were extended to $T=500$ (Figure 3(b), Figure 4). After the initial growth phase, the simulation remains relatively stable through the entire nonlinear phase. Furthermore, we observe clear noise reduction in the nonlinear phase when applying the pseudoinverse resampling algorithm. In fact, while dynamics vary between runs, every run with resampling shows complete noise reduction throughout the full simulation. The ability of the resampled simulation to remain stable and preserve coherent phase-space structures over many plasma periods provides strong evidence that the PIC framework is correctly capturing both the linear and nonlinear dynamics of the two-stream instability and is not dominated by numerical noise, grid artifacts, or energy drift.
With resampling enabled, however, we observe small but regular oscillations in the electric field strength during the nonlinear phase and after the initial vortices merge. As shown in Fig. 4, secondary vortices form and interact after $T=32$, trapping particles whose oscillatory motion in $x$ modulates the charge density and electric field strength. This may be a signature of nonlinear bounce dynamics, common in two-stream instability tests. Alternatively, these oscillations may arise from numerical instabilities, similar to those discussed in 7.1. Fine-scale structures developed at later times will produce grid-scale oscillations which can alias onto the physical modes modulating the electric field. Both of these, physical and numerical, effects can appear similarly in the results and are therefore a topic for further study in future work.
@@ 8 Side effects of resampling
Figure 7(a) demonstrates that the dynamics of the plasma is not highly sensitive to the resampling period within a broad range (e.g., 181-8 $\omega_{p}^{-1}]$), but that rebound starts earlier with higher resampling rates, indicating that resampling is disturbing the plasma to some extent. To investigate this effect, Figure 12 (left) shows a linear example with resampling every time step and increasing resampling periods. The highest resampling rates clearly lead to instabilities and these instabilities demonstrate classic growing oscillations up to a small amplitude, see detail in Figure 12 (right), that is growing in mean value and plateaus at about $1.0$. Perhaps stability analysis could provide insight on this issue.
@@
While this instability at high resampling rates is not a practical problem, per se, in that all of our tests are stabilized with much lower resampling frequencies, this data does indicate that resampling can have adverse effects. This test problem is highly idealized and does not provide convincing experimental evidence that the adaptivity strategy of simply mapping to the original grid will be robust in practice. More sophisticated adaptivity strategies are the subject of future work and would require more scientifically relevant models such as problems in magnetic reconnection in at least $2X+3V$.
@@ 8.1 Discussion
Figure 10 sheds light onto the complex relationship and dependency between all the simulation variables, i.e., resampling rate, spatial and velocity grid density and time step. It is difficult to draw definitive conclusions about what exactly is driving the dynamics of the electric field damping and rebounding. As discussed in §7, the grid resolution plays a large role in reducing the interaction of subgrid modes with the physical waves, while the particle density reduces stochastic, noise-driven heating effects in the plasma. These effects do not, however, converge monotonically suggesting that they are linked in some way. This requires further study in future work.
The inclusion of collisions may also provide a smooth mechanism. The collision operator acts as a controlled, physically motivated diffusion term in the kinetic equations. This diffusive term continuously smooths out any build up of high-frequency components in the distribution function which can lead to spurious rebounds. In other algorithms (Fig. 5.5 [KrausThesis]), clean decays to machine precision have been shown using collisions. Collisions must be used with some level of caution, however. Much of the essential collisionless interactions in the Landau damping test, in particular in the nonlinear regime, rely on fine-scale interactions between the particles and waves. Introducing a collision operator can have the effect of smoothing out those essential fine-scale interactions which, if too strong, will disrupt the fundamental physics of the system. Recent work in the PETSc library [AdamsHirvijokiKnepleyBrownIsaacMills2017, pusztay2023landau, Adams2022a, adams2024performance] has provided a variety of collision algorithms that can be included in the PETSc-PIC framework in future work.
@@ 9 Conclusion
The paper develops a new approach to particle resampling that uses a conservative projection, a pseudoinverse, to map any distribution of particles to essentially any other distribution of particle while conserving all moments up to the degree of polynomial that the projection function space can represent exactly. This method is evaluated with a static particle and continuum Cartesian grids, simply remapping to the original particle grid, on standard two-stream instability and Landau damping test problems. Two-stream instability tests show minor accuracy improvements when using resampling methods, as well as a well maintained particle grid over long time simulations. In Landau damping tests, where noise is problematic, the linear cases, show that resampling reduces noise considerably and coherent dynamics are maintained for long times, where as the solution becomes essentially all noise without resampling.
@@
With entropy measures, from our particle Landau collision operator [pusztay2023landau], we can determine the continuum grids required for resampling or a continuum collision operator [adams2024performance, pusztay2023landau] to keep entropy generation by the projection well below the entropy generated by the physics.
@@
AMR in velocity and regular particle grids on each cell of an adapted grid, like a cubed sphere [adams2024performance], is a path for generating adapted particle grids through continuum mesh adaptation.
@@ Appendix A Artifact description and reproducibility
PETSc output files with all data, provenance information, and reproducibility instructions for all tables and plots can be obtained from git@gitlab.com:markadams4/resampling-paper.git. This includes the python scripts that generates the plots and run scripts, makefiles and PETSc resource files used to generate the data, and the test harness code in src. The src/A.X directories has data for A=0.A=0.X. The exact PETSc versions (SHA1) are in the data files, with the provenance data, all parameters used in each test, but any PETSc version from v3.22 should suffice to reproduce this data.